OpenStudy (anonymous):

A sequence has its first term equal to 4, and each term of the sequence is obtained by adding 2 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 2 and f(n) = f(n - 1) + 4; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2n; n > 1 f(1) = 2 and f(n) = f(n - 1) + 4n; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2; n > 1

OpenStudy (cwrw238):

well the first term is 4 so which 2 options can you rule out?

OpenStudy (anonymous):

A and C?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

now try which of the other 2 options add 2 to a term in the series/

OpenStudy (cwrw238):

try plugging in 2 for n and see if it gives 6 as the second term (we know that the series is 4,6,8 ...)

OpenStudy (anonymous):

B?

OpenStudy (cwrw238):

we know its either b or d

OpenStudy (anonymous):

i think its B

OpenStudy (cwrw238):

lets try b: f(n-1) + 2n f(2) = f(2-1) + 2(2) = f(1) + 4 now f(1) = 4 so f(20 = 4+4 = 8 this one?

OpenStudy (cwrw238):

do you follow theabove?

OpenStudy (anonymous):

yea but did you mean to put(2)

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

if you try d you get f(1) + 2 = 4 + 2 = 6

OpenStudy (anonymous):

sorry i was talking to my teacher but so its D?

OpenStudy (cwrw238):

yes